If X is normally distributed with mean 150.3cm and variance 25. Find the probability that X is more than 10cm difference from the mean


1
Expert's answer
2021-03-19T11:45:09-0400

"X\sim N(\mu, \sigma^2 )"

Given "\mu=150.3, \sigma^2=25"


"P(X<140.3)=P(Z<\dfrac{140.3-150.3}{\sqrt{25}})"

"=P(Z<-2)\approx0.022750"


"P(X>160.3)=1-P(X\leq 160.3)"

"=1-P(Z\leq \dfrac{160.3-150.3}{\sqrt{25}})=1-P(Z\leq 2)"

"\approx0.022750"

"P(difference>10)"

"=P(X<140.30)+P(X>160.3)"

"\approx0.022750+0.022750=0.0455"

Check

"\sigma=\sqrt{25}=5"

"10=2\sigma"

By the 68–95–99.7 rule


"P(\mu-2\sigma\leq X\leq \mu+2\sigma)\approx0.9545"

"P(difference>10)\approx1-0.9545=0.0455"